Use the following information to determine tan(2x). sin(x) 2 and cos(x) is positive y6
Answer #1

Slution Sind= -2 6 al Cosa es Positive We know that Sinal + 65 x = 1 (12) + Case=1 4 + @s?d=1 Cosx = 2 6 Cosx = 1/2 Corn= t1this tam (2x) = 2tona 1- tamne (2)(-5) 1- (-л) - 2 1 - 2 taman = ar Ansues.